# Reward-Augmented Data Enhances Direct Preference Alignment of LLMs

### v2

@@ 1 Introduction
To address these issues, we propose learning reward-conditioned policies as a straightforward fix to the above issues. By optimizing the LLM to generate responses conditioning on their qualities, the model is allowed to discern and leverage patterns within responses of varied quality. As a result, learning from both chosen and rejected responses alleviates the issue of unnecessarily unlearning high-quality rejected responses; distinguishing between varying-quality chosen responses alleviates the issue of indiscriminately accepting low-quality ones. By identifying common patterns in responses of similar quality and distinguishing them from those of differing quality, the LLM becomes more adept at generalizing to more optimal responses that are sparse in data.
@@ 2 Background
The RLAIF framework involves two major steps: preference dataset construction with AI feedback and preference optimization. As a surrogate for human preference, AI feedback, including LLM-as-Judge (Zheng et al., 2024; Cui et al., 2023) and Reward-Model-as-Judge (Adler et al., 2024; Dong et al., 2024), can be used to rank responses and generate preference pairs. Specifically, consider the judge model $r(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ that outputs a scalar reward value representing the quality of $y$ under $x$. For each prompt $x\in\mathcal{X}$, two responses, $y_{1}$ and $y_{2}$, are independently sampled—either from the same reference model (Xiong et al., 2024; Wu et al., 2024) or several different models (Zhu et al., 2023; Zhang et al., 2024). Then $r(x,y_{1})$ and $r(x,y_{2})$ are evaluated to determine the preferred response $y_{w}=\mathop{\mathrm{argmax}}_{y\in\{y_{1},y_{2}\}}r(x,y)$ and dispreferred response $y_{l}=\mathop{\mathrm{argmin}}_{y\in\{y_{1},y_{2}\}}r(x,y)$. By sampling responses and ranking them for a set of $N$ prompts, we get a preference dataset: $\mathcal{D}_{N}=\{x^{i},y_{w}^{i},y_{l}^{i},r(x^{i},y_{w}^{i}),r(x^{i},y_{l}^{i})\}_{i=1}^{N}$.
@@
$\mathcal{L}_{\text{DPO}}(\pi;\mathcal{D}_{N})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\biggl{[}\log\sigma\biggl{(}\beta\log\frac{\pi(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\biggr{)}\biggr{]},$
@@ 3.1 Limitations of Direct Alignment from Preference
The dataset $\mathcal{D}_{N}$ often contains preference pairs where the rejected response $y_{l}$ is only marginally worse than the chosen one $y_{w}$. Direct alignment algorithms, however, primarily focus on relative preferences and are unaware of the responses’ quality values and gaps. Striving to maximize the reparameterized reward gap between the chosen and rejected responses will risk overfitting and the unnecessary unlearning of high-quality responses, potentially diminishing the model’s performance by discarding valuable alternatives. Furthermore, in such a finite data regime where only a sample estimate of the true preference is accessible, it can be very possible that $p^{*}(y_{l}\succ y_{w}\mid x)>0.5$, i.e., $y_{l}$ is in fact more preferred than $y_{w}$ in expectation. This issue becomes even more pronounced when the preference data generated with the imperfect judge model is noisy.
We illustrate this limitation with the example in Table 2. For $\mathcal{D}_{N=1}$ that contains a single preference pair [^1] For simplicity, we write $(x,y_{w},y_{l},r(x,y_{w}),r(x,y_{l}))\in\mathcal{D}_{N}$ as $y_{w}>y_{l}$. with reward $r(x,y_{1})=9$ and $r(x,y_{2})=8$, the optimal policy learned from $\mathcal{D}_{N=1}$ is $\pi^{*}(y_{1}\mid x)=1$. This causes the model to avoid generating $y_{2}$, a response of nearly equivalent quality.
@@
Preference data often contains responses that, despite being preferred in pairwise comparisons, exhibit substantial variation in quality. As a result, the optimal responses—those associated with the highest reward values—are sparse in the dataset. Since direct alignment algorithms do not account for these reward values, the trained model struggles to differentiate between responses of varying quality and fails to generalize effectively to the sparse optimal responses.
@@
Table 1: High-quality rejected responses such as $y_{2}$ can be unnecessarily unlearned: $\pi^{*}(\cdot\mid x)$ deterministically generates $y_{1}$.
@@
3.2 Reward-Conditioned Policies Resolve the Limitations
A straightforward way to address the limitations of direct alignment algorithms—specifically, their inability to account for the quality of responses—is to optimize a reward-conditioned policy. In this approach, the LLM policy is trained to generate responses corresponding to different reward values, enabling it to become aware of and adapt to these reward distinctions. By doing so, the LLM not only learns the patterns associated with the preferred responses but also retains the valuable information from the rejected ones, preventing the unlearning of high-quality rejected responses. For example, in Table 2, reward-conditioned policies learn to generate both $y_{1}$ and $y_{2}$, instead of unlearning $y_{2}$. This reward-based conditioning also enhances the model’s ability to differentiate between responses of varying quality, even if both are preferred over a rejected alternative, as illustrated in Table 2. Besides, by extracting common patterns across responses with different quality levels, the LLM becomes more generalizable and is capable of generating the highest-quality responses (e.g., $r=10$), which are often sparse in the training preference data.
@@ 4 Method
With the above motivation, we propose a data relabeling method that constructs a reward-augmented dataset by conditioning the preference pairs on the reward values given by the judge model. Specifically, we define the goal-conditioned reward $r_{g}(x,y)=\mathmybb{1}\bigl{(}g=r(x,y)\bigr{)}$ as a function of the quality reward $r$. The objective of the reward-conditioned policy $\pi(y\mid x;g)$ is thus to minimize the absolute difference between the goal reward $g$ and the response reward $r(x,y)$, which is equivalent to maximizing the goal-conditioned reward $r_{g}(x,y)$, i.e.,
$\min_{\pi}\mathbb{E}_{g,x\sim\mathcal{D}_{N},y\sim\pi(\cdot\mid x;g)}\bigl{[}|r(x,y)-g|\bigr{]}=\max_{\pi}\mathbb{E}_{g,x\sim\mathcal{D},y\sim\pi(\cdot\mid x;g)}\bigl{[}r_{g}(x,y)\bigr{]}.$ (4.1)
To optimize the RHS of Equation (4.1), we first observe that under the new goal-conditioned reward metric $r_{g}$, for each preference pair ${x^{i},y_{w}^{i},r_{w}^{i},y_{l}^{i},r_{l}^{i}}$ in $\mathcal{D}_{N}$, we have
$r_{g=r_{w}^{i}}(x,y_{w}^{i})=1,r_{g=r_{w}^{i}}(x,y_{l}^{i})=0,r_{g=r_{l}^{i}}(x,y_{l}^{i})=1,r_{g=r_{l}^{i}}(x,y_{w}^{i})=1.$
@@
We illustrate our method in Figure 1. For each preference pair with index $i$ in $\mathcal{D}_{N}$, two goals are defined, corresponding to the reward values of the chosen response $y_{w}^{i}$ and the rejected response $y_{l}^{i}$. Specifically, under the first goal $g=y_{w}^{i}$, the relabeled rewards are $r_{g}(x,y_{w}^{i})=1$ and $r_{g}(x,y_{l}^{i})=0$. The original ranking of responses remains the same, except that the LLM is preference optimized conditioned on $g=y_{w}^{i}$. Similarly, under the second goal $g=y_{l}^{i}$, the relabeled rewards are $r_{g}(x,y_{l}^{i})=1$ and $r_{g}(x,y_{w}^{i})=0$. Thus, the chosen and rejected responses are reversed as $y_{l}^{i}$ and $y_{w}^{i}$, respectively. By generating preference pairs conditioned on the goal reward for both the chosen and rejected responses, we obtain a reward-augmented dataset of size $2N$. Finally, this new dataset can be used with any direct alignment algorithm, such as DPO.
In this work, we implement the reward-conditioned policy $\pi(y\mid x;g)$ as the LLM with a system prompt (or a prefix before the user prompt $x$ if system prompts are not supported by the LLM) such as “generate responses of score $g$”. At inference time, the LLM is conditioned on the highest possible reward value, e.g., $g=10$, to generate the responses.
@@ 5 Related Work
In order for the LLMs to follow instructions and better align with human intents, it is common practice to build a preference dataset containing a set of prompts and a pair of responses for each prompt, whose qualities are ranked by humans (Ouyang et al., 2022) or judge models (Bai et al., 2022b). A popular pipeline (Cui et al., 2023; Tunstall et al., 2023b; Wang et al., 2024c; Ivison et al., 2023; Zhu et al., 2023) for constructing offline (i.e., fixed) datasets involves sampling off-policy responses from various LLMs for each prompt in the hope to increase the response diversity. The preference data can also be generated online (Guo et al., 2024) or iteratively (Bai et al., 2022a; Xu et al., 2023; Gulcehre et al., 2023; Hoang Tran, 2024; Xiong et al., 2023; Dong et al., 2024; Calandriello et al., 2024; Rosset et al., 2024) by sampling and ranking on-policy responses from the training LLM. Recent works (Zhang et al., 2024; Cen et al., 2024; Xie et al., 2024) have also proposed systematically exploring the responses online and actively eliciting the preference. The proposed method in this paper is orthogonal to the construction ways of the preference data and can be applied to any dataset created either off-policy or on-policy.
@@
Conditioning LLMs during training has proven effective for aligning responses with specific human objectives. SteerLM (Dong et al., 2023b; Wang et al., 2023b) extends SFT by conditioning the LLM on the multi-dimensional annotated attributes in data, such as humor and toxicity, in order to steer model responses with user customizability. Directional Preference Alignment (DPA) (Wang et al., 2024a) proposed a variant of rejection sampling fine-tuning (Yuan et al., 2023; Dong et al., 2023a) that conditions on the direction of the multi-objective reward, i.e., a user-dependent linear combination of the reward attributes (helpfulness and verbosity in their experiments), that represents diverse preference objectives. These methods aim to train a single LLM that can flexibly adjust to various user preference profiles. On the contrary, our method targets the limitations of direct alignment algorithms by introducing reward-augmented relabeling. This also differs from Conditioned-RLFT (Wang et al., 2023a), which leverages the data source information by learning a class-conditioned policy with RL-free supervised learning. Reward-aware Preference Optimization (RPO), introduced in Nemotron-4 (Adler et al., 2024), attempts to approximate the reward gap using the implicit reward and is motivated to resolve the unlearning issues of DPO, which our work also addresses. However, we show that more limitations beyond unlearning can be simply fixed with reward-conditioned LLMs and propose an easy-to-implement data relabeling method that integrates seamlessly with any direct alignment algorithm. We compare with all the aforementioned methods in our experiments.
@@ 6.1 Reward-Augmented Data Boosts DPO Performance
We begin by conducting experiments to demonstrate that applying the proposed method to fixed offline preference datasets leads to consistent performance improvements in DPO.

### v3

@@ 1 Introduction
To address these issues, we propose learning reward-conditioned policies as a straightforward fix. By optimizing the LLM to generate responses conditioning on their qualities, the model is allowed to discern and leverage patterns within responses of varied quality. As a result, learning from both chosen and rejected responses alleviates the issue of unnecessarily unlearning high-quality rejected responses; distinguishing between varying-quality chosen responses alleviates the issue of indiscriminately accepting low-quality ones. By identifying common patterns in responses of similar quality and distinguishing them from those of differing quality, the LLM becomes more adept at generalizing to more optimal responses that are sparse in data.
@@ 2 Background
The RLAIF framework involves two major steps: preference dataset construction with AI feedback and preference optimization. As a surrogate for human preference, AI feedback, including LLM-as-Judge (Zheng et al., 2024; Cui et al., 2023) and Reward-Model-as-Judge (Adler et al., 2024; Dong et al., 2024), can be used to rank responses and generate preference pairs. Specifically, consider the judge model $r(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ that outputs a scalar reward value representing the quality of $y$ under $x$. For each prompt $x\in\mathcal{X}$, two responses, $y_{1}$ and $y_{2}$, are independently sampled—either from the same reference model (Xiong et al., 2024; Wu et al., 2024) or several different models (Zhu et al., 2023a; Zhang et al., 2024). Then $r(x,y_{1})$ and $r(x,y_{2})$ are evaluated to determine the preferred response $y_{w}=\mathop{\mathrm{argmax}}_{y\in\{y_{1},y_{2}\}}r(x,y)$ and dispreferred response $y_{l}=\mathop{\mathrm{argmin}}_{y\in\{y_{1},y_{2}\}}r(x,y)$. By sampling responses and ranking them for a set of $N$ prompts, we get a preference dataset: $\mathcal{D}_{N}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$. For the simplicity of our discussions, we assume that the reward function $r$ is bounded in $[0,r_{\text{max}}]$.
@@
$\mathcal{L}_{\text{DPO}}(\pi;\mathcal{D}_{N})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{N}}\biggl{[}\log\sigma\biggl{(}\beta\log\frac{\pi(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\biggr{)}\biggr{]},$
@@ 3.1 Limitations of Direct Alignment from Preference
The dataset $\mathcal{D}_{N}$ often contains preference pairs where the rejected response $y_{l}$ is only marginally worse than the chosen one $y_{w}$. Direct alignment algorithms, however, primarily focus on relative preferences and are unaware of the responses’ quality values and gaps. Striving to maximize the reparameterized reward gap between the chosen and rejected responses will risk overfitting and unnecessary “unlearning”, i.e., probability decrease, of high-quality responses, potentially diminishing the model’s performance by discarding valuable alternatives. Furthermore, in such a finite data regime where only a sample estimate of the true preference is accessible, it can be very possible that $p^{*}(y_{l}\succ y_{w}\mid x)>0.5$, i.e., $y_{l}$ is in fact more preferred than $y_{w}$ in expectation. This issue becomes even more pronounced when the preference data generated with the imperfect judge model is noisy.
We illustrate this limitation with the example in Table 2, where we define the maximum reward $r_{\text{max}}$ as $10$. For $\mathcal{D}_{N=1}$ that contains a single preference pair [^1] For simplicity, we write $(x,y_{w},y_{l})\in\mathcal{D}_{N}$ as $y_{w}\succ y_{l}$. with reward $r(x,y_{1})=9$ and $r(x,y_{2})=8$, the optimal policy learned from $\mathcal{D}_{N=1}$ is $\pi^{*}(y_{1}\mid x)=1$. This causes the model to avoid generating $y_{2}$, a response of nearly equivalent quality.
@@
Preference data often contains responses that, despite being preferred in pairwise comparisons, exhibit substantial variation in quality. As a result, the optimal responses—those associated with the highest reward value $r_{\text{max}}$—are sparse in the dataset. Since direct alignment algorithms do not account for these reward values, the trained model struggles to differentiate between responses of varying quality and fails to generalize effectively to the sparse optimal responses.
@@
$\pi^{*}(y\mid x,g=9)$ 1 0
$\pi^{*}(y\mid x,g=8)$ 0 1
Table 1: High-quality rejected responses such as $y_{2}$ can be unnecessarily unlearned: $\pi^{*}(\cdot\mid x)$ deterministically generates $y_{1}$. Reward-conditioned policies learn both responses and are easier to generalize to $g=10$ with the extracted features from $g=8$ and $g=9$.
@@
3.2 Reward-Conditioned Policies Learn from the Full Spectrum of Response Quality
A straightforward way to address the limitations of direct alignment algorithms—specifically, their inability to account for the quality of responses—is to optimize a reward-conditioned policy. In this approach, the LLM policy is trained to generate responses corresponding to different reward values, enabling it to become aware of and adapt to these reward distinctions. By doing so, the LLM not only learns the patterns associated with the preferred responses but also retains the valuable information from the rejected ones, preventing the unlearning of high-quality rejected responses. For example, in Table 2, reward-conditioned policies learn to generate both $y_{1}$ and $y_{2}$, instead of unlearning $y_{2}$. This reward-based conditioning also enhances the model’s ability to differentiate between responses of varying quality, even if both are preferred over a rejected alternative, as illustrated in Table 2. Besides, by extracting common patterns across responses with different quality levels, the LLM becomes more generalizable and is capable of generating the highest-quality responses with reward $r_{\text{max}}$ (e.g., $r_{\text{max}}=10$), which are often sparse in the training preference data.
@@ 4 Method
With the above motivation, we propose a data relabeling method that constructs a reward-augmented dataset by conditioning the preference pairs on the reward values given by the judge model $r$. Specifically, we define the goal-conditioned reward function $R(x,y,g)=-(g-r(x,y))^{2}$ as a function of the reward function $r$. The objective of the reward-conditioned policy $\pi(y\mid x,g)$ is thus to minimize the square difference between the goal reward $g$ and the response reward $r(x,y)$, which is equivalent to maximizing the goal-conditioned reward $R(x,y,g)$, i.e.,
$\min_{\pi}\mathbb{E}_{g,x\sim\mathcal{D}_{N},y\sim\pi(\cdot\mid x,g)}\bigl{[}(g-r(x,y))^{2}\bigr{]}=\max_{\pi}\mathbb{E}_{g,x\sim\mathcal{D},y\sim\pi(\cdot\mid x,g)}\bigl{[}R(x,y,g)\bigr{]}.$ (4.1)
To optimize the RHS of Equation (4.1), we first observe that under the new goal-conditioned reward metric $r$, for each preference pair ${x^{i},y_{w}^{i},r_{w}^{i},y_{l}^{i},r_{l}^{i}}$ in $\mathcal{D}_{N}$, we have
$R(x,y_{w}^{i},g=r_{w}^{i})$ $=0>R(x,y_{l}^{i},g=r_{w}^{i})=-(r_{w}^{i}-r_{l}^{i})^{2},$
$R(x,y_{l}^{i},g=r_{l}^{i})$ $=0>R(x,y_{w}^{i},g=r_{l}^{i})=-(r_{w}^{i}-r_{l}^{i})^{2}.$
@@
We illustrate our method in Figure 1. For each preference pair with index $i$ in $\mathcal{D}_{N}$, two goals are defined, corresponding to the reward values of the chosen response $y_{w}^{i}$ and the rejected response $y_{l}^{i}$. Specifically, under the first goal $g=r_{w}^{i}$, the relabeled rewards are $R(x,y_{w}^{i},g)=0$ and $R(x,y_{l}^{i},g)=-(r_{w}^{i}-r_{l}^{i})^{2}$. The original ranking of responses remains the same, except that the LLM is preference optimized conditioned on $g=r_{w}^{i}$. Similarly, under the second goal $g=r_{l}^{i}$, the relabeled rewards are $R(x,y_{l}^{i},g)=0$ and $R(x,y_{w}^{i},g)=-(r_{w}^{i}-r_{l}^{i})^{2}$. Thus, the chosen and rejected responses are reversed as $y_{l}^{i}$ and $y_{w}^{i}$, respectively. By generating preference pairs conditioned on the goal reward for both the chosen and rejected responses, we obtain a reward-augmented dataset of size $2N$. Finally, this new dataset can be used with any direct alignment algorithm, such as DPO.
In this work, we implement the reward-conditioned policy $\pi(y\mid x,g)$ as the LLM with a system prompt (or a prefix before the user prompt $x$ if system prompts are not supported by the LLM) such as “generate responses of score $g$”. At inference time, the LLM is conditioned on the optimal goal $g^{\star}=r_{\text{max}}$ that is the highest possible reward value, e.g., $g^{\star}=r_{\text{max}}=10$, to generate the responses.
We provide the following theoretical guarantees for our method (see A.4 for a formal description).
Theorem 4.1 (Informal).
Let $J(\pi)=\mathbb{E}_{x\sim d_{0},y\sim\pi(\cdot|x,g^{\star})}\big{[}R(x,y,g^{\star})\big{]}$ be the performance measure, where $R$ denotes the ground-truth goal-conditioned reward function and $g^{\star}$ denotes the optimal goal. Under mild assumptions, the policy $\widehat{\pi}$ optimized from the reward-augmented DPO with a SFT regularizer satisfies that with probability at least $1-\delta$,
$J(\pi^{*})-J(\widehat{\pi})\!\leq\!\sqrt{\frac{1}{N}}\cdot\bigg{\{}\frac{\sqrt{6}}{4}\big{(}1+\exp(B)\big{)}^{2}\big{(}\big{(}C_{\mu_{\bar{\mathcal{D}}}}(\mathcal{R};\pi^{\star},\pi_{\text{sft}})\big{)}^{2}+1\big{)}\iota$
$\qquad\qquad\qquad+\mathbb{E}_{x\sim d_{0}}\Big{[}\mathrm{KL}\big{(}\pi^{\star}(\cdot|x,g^{\star})\|\pi_{\mathrm{ref}}(\cdot|x,g^{\star})\big{)}\Big{]}\bigg{\}},$ (4.2)
where $\pi^{*}=\mathop{\mathrm{argmax}}_{\pi}J(\pi)$ and $\iota=\sqrt{\log\left(\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})/\delta\right)}$ with $\varepsilon=(6\cdot(1+e^{B})\cdot N)^{-1}$. Here, $N$ denotes the number of preference pairs in $\mathcal{D}$, $B$ denotes the upper bound of the reward models, and the partial coverage coefficient $C_{\mu_{\bar{\mathcal{D}}}}(\mathcal{R};\pi^{\star},\pi_{\text{sft}})$ is defined in Assumption A.3.
The detailed proof is provided in A.5. The above theorem shows that our method attains global convergence to the optimal policy and the suboptimality decays at the order of $N^{-1/2}$ ($N$ denotes the size of the reward-augmented preference dataset), which provides a theoretical justification for the strong empirical performance of the introduced reward-augmented DPO. Unlike prior works on goal-conditioned RL with supervised learning (Yang et al., 2022; Ghosh et al., 2019), which typically establish weaker results such as local performance improvements or the optimization of a lower bound on $J(\pi)$, our analysis guarantees global convergence to the optimal policy. This distinction underscores the significance of integrating DPO-like methods with goal-conditioned approaches.
@@ 5 Related Work
In order for the LLMs to follow instructions and better align with human intents, it is common practice to build a preference dataset containing a set of prompts and a pair of responses for each prompt, whose qualities are ranked by humans (Ouyang et al., 2022) or judge models (Bai et al., 2022b). A popular pipeline (Cui et al., 2023; Tunstall et al., 2023b; Wang et al., 2024c; Ivison et al., 2023; Zhu et al., 2023a) for constructing offline (i.e., fixed) datasets involves sampling off-policy responses from various LLMs for each prompt in the hope to increase the response diversity. The preference data can also be generated online (Guo et al., 2024) or iteratively (Bai et al., 2022a; Xu et al., 2023; Gulcehre et al., 2023; Hoang Tran, 2024; Xiong et al., 2023; Dong et al., 2024; Calandriello et al., 2024; Rosset et al., 2024) by sampling and ranking on-policy responses from the training LLM. Recent works (Zhang et al., 2024; Cen et al., 2024; Xie et al., 2024) have also proposed systematically exploring the responses online and actively eliciting the preference. The proposed method in this paper is orthogonal to the construction ways of the preference data and can be applied to any dataset created either off-policy or on-policy.
@@
Conditioning LLMs during training has proven effective for aligning responses with specific human objectives. SteerLM (Dong et al., 2023b; Wang et al., 2023b) extends SFT by conditioning the LLM on the multi-dimensional annotated attributes in data, such as humor and toxicity, in order to steer model responses with user customizability. Directional Preference Alignment (DPA) (Wang et al., 2024a) proposed a variant of rejection sampling fine-tuning (Yuan et al., 2023; Dong et al., 2023a) that conditions on the direction of the multi-objective reward, i.e., a user-dependent linear combination of the reward attributes (helpfulness and verbosity in their experiments), that represents diverse preference objectives. These methods aim to train a single LLM that can flexibly adjust to various user preference profiles. On the contrary, our method targets the limitations of direct alignment algorithms by introducing reward-augmented relabeling. This also differs from Conditioned-RLFT (Wang et al., 2023a), which leverages the data source information by learning a class-conditioned policy with RL-free supervised learning. Reward-aware Preference Optimization (RPO), introduced in Nemotron-4 (Adler et al., 2024), attempts to approximate the reward gap using the implicit reward and is motivated to resolve the unlearning issues of DPO, which our work also addresses. However, we show that more limitations beyond unlearning can be simply fixed with reward-conditioned LLMs and propose an easy-to-implement data relabeling method that integrates seamlessly with any direct alignment algorithm.
@@ 6.1 Reward-Augmented Data Boosts DPO Performance
We begin by conducting experiments to demonstrate that applying the proposed method on DPO to fixed offline preference datasets leads to consistent performance improvements.
Setup.
We adopt the UltraFeedback (Cui et al., 2023) preference dataset containing reward values scored by GPT-4 (LLM-as-Judge) that is ranged between $1$ and $10$ for each of the preference pairs. Our method constructs reward-augmented data by conditioning on these judge values. We fine-tune on various open-weight LLMs, including Mistral-7B-Instruct-v0.3, Qwen2-7B-Instruct, Llama-3.1-8B-Instruct, Gemma-2-9B-It, and SPPO (fine-tuned from Gemma2-9B-It). We use the DPO implementation in the Huggingface Alignment Handbook (Tunstall et al., 2023a). The hyperparameters and prompts that we use are listed in Appendix B.1.
Results.